OpenStudy (anonymous):

simplify -2x+4+4x=5x+3+x

OpenStudy (anonymous):

4x=1

OpenStudy (anonymous):

x=1/4

jimthompson5910 (jim_thompson5910):

-2x+4+4x=5x+3+x 2x+4=5x+3+x 2x+4=6x+3 2x=6x+3-4 2x-6x=3-4 -4x=3-4 -4x=-1 x=(-1)/(-4) x=1/4

OpenStudy (anonymous):

-2x+4+4x=5x+3+x Since -2x and 4x are like terms, subtract 4x from -2x to get 2x. 2x+4=5x+3+x Since 5x and x are like terms, add x to 5x to get 6x. 2x+4=6x+3 Since 6x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 6x from both sides. 2x+4-6x=3 Since 2x and -6x are like terms, add -6x to 2x to get -4x. -4x+4=3 Since 4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4 from both sides. -4x=-4+3 Add 3 to -4 to get -1. -4x=-1 Divide each term in the equation by -4. -(4x)/(-4)=-(1)/(-4) Simplify the left-hand side of the equation by canceling the common terms. x=-(1)/(-4) Simplify the right-hand side of the equation by simplifying each term. x=(1)/(4)